# JWST’s First View of Tidal Disruption Events: Compact, Accretion-Driven Emission Lines & Strong Silicate Emission in an Infrared-selected Sample

### v1

@@ 5 Discussion & Conclusion
Motivated by the strong silicate emission features, we have modeled the JWST MIRI spectra with a simple, optically thin dust model that accounts for the expanding isodelay surface. This model can accurately reproduce the strong silicate features and the long wavelength continuum emission. However, the current model assumes a $\delta$-function for the input flare, while TDEs are observed to flare for $\sim$100s of days (see van Velzen et al., 2020; and references therein) and most even show years-long optical and UV plateaus associated with the viscous spreading of a newly formed accretion disk (Mummery et al., 2024). The late-time emission from the TDE will heat the surviving dust near the SMBH, but since the initial flare is quite luminous, most of the small grains and silicates at such small radii will have been destroyed. Figure 6 shows that dust emission from larger ($a=0.3$ µm), pure graphite grains heated by a late-time plateau of $L_{\mathrm{late-time}}\approx L_{\mathrm{init}}/25$ can reproduce the near-IR excess at $\lambda\lesssim 8$ µm. Interestingly, near-IR excesses are also seen in AGN and have been similarly explained with hot ($T\gtrsim 1200$ K), pure graphite dust that can exist closer to the SMBH than silicate dust (e.g. Mor et al., 2009; Mor & Netzer, 2012). As TDEs may emit a sizable fraction of their total energy at late times in a viscously spreading accretion disk (Mummery & Balbus, 2020; Mummery, 2021; Mockler & Ramirez-Ruiz, 2021; Mummery & van Velzen, 2024), accurate estimation of the late-time plateaus and by association the close dust to the SMBH is crucial to our understanding of TDE energetics and solving the TDE missing energy problem (e.g. Lu & Kumar, 2018). We stress, however, that the inclusion of the full TDE light curve in our modeling may change these constraints (e.g. to smaller $a$ or lower $L_{\mathrm{late-time}}$), as the response of dust at intermediate radii may produce significant emission in the 5-8 µm  regime.
@@
Given the complexity associated with astrophysical dust in extreme environments, our modeling has made a number of key assumptions and approximations. In addition to the $\delta$-function model for the intrinsic flare, another important assumption that we have made is that the dust is optically thin to its own emission (i.e. $\tau_{9.7\,\mu\mathrm{m}}\lesssim 1$). This assumption allows us to neglect scattering in our calculations, as a proper treatment of scattering is difficult to do in a non-spherical emitting surface given geometric effects. Our order of magnitude estimates suggest that optically thin dust is relevant for the timescales (i.e. distances) we are probing with JWST MIRI. That said, dense, optically thick dust closer to the SMBH may impact our calculations as it will attenuate the dominant UV emission and lead to the dust being heated by its own reprocessed radiation, rather than the intrinsic flare. Future efforts with Monte Carlo radiation transport simulations are necessary to assess the impact of scattering and heating from reprocessed emission on with the necessary time-dependent effects that exist in TDEs. Such simulations are now starting to be viable computationally, but to-date have only been done with gray opacities to estimate the light curve evolution (e.g. Tuna et al., 2025).
@@
This modeling makes two key predictions that are relevant for next-generation IR observatories. At very late times ($\gtrsim 100$ years), the illumination of distant, cool dust will lead to far-IR dust echoes that may be detectable with next-generation far-IR observatories like the PRobe far-Infrared Mission for Astrophysics (PRIMA; Moullet et al., 2023). At very early times ($\lesssim 0.5$ years), the spectrum should be dominated by pure graphite dust, as all of the silicates near the SMBH will be destroyed. JWST observations in the first 6 months after a TDE are necessary to test this hypothesis and inform our modeling of the dust close to the SMBH. Although MIR-selected TDE candidate samples are contaminated by SNe at early times, recent work has shown that only a small fraction ($\sim 0.1\%$) of the dominant contaminants, Type Ia SNe, are MIR-bright (Mo et al., 2025), thereby making these early-time searches more tractable. Additionally, in the coming years, numerous new observatories will revolutionize time-domain IR studies with greatly improved cadence, including the Nancy Grace Roman Space Telescope (Spergel et al., 2015), the Wide-field INfrared Transient ExploreR (WINTER; Lourie et al., 2020), and the PRime-focus Infrared Microlensing Experiment (PRIME [^3] http://www-ir.ess.sci.osaka-u.ac.jp/prime/index.html) in the near-IR and NEOSurveyor (Mainzer et al., 2023) in the mid-IR. Likewise, SPHEREx (Doré et al., 2014), which is expected to launch in March 2025, will produce all-sky near-to-mid-IR spectra every 6 months that can help identify IR-bright TDEs earlier in their evolution. Thus, we expect that catching TDE dust echoes in their early phases of evolution will be possible with the next generation missions. These observations will break degeneracies between dust composition and accretion luminosity, thereby allowing us to answer fundamental questions related to the overall impact of TDEs on black hole growth.
@@
This work is based on observations made with the NASA/ESA/CSA JWST. The data were obtained from the Mikulski Archive for Space Telescopes at the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-03127 for JWST. These observations are associated with program #3696. Support for program #3696 was provided by NASA through a grant from the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-03127. We acknowledge the support of the National Aeronautics and Space Administration through ADAP grant number 80NSSC24K0663. MG is supported in part by NASA XMM-Newton grant 80NSSC24K1885.

### v2

@@ 5 Discussion & Conclusion
The exquisite resolving power and angular resolution of JWST MIRI/MRS has allowed us to disentangle the MIR nuclear emission for four nearby ($D_{L}\lesssim 200$ Mpc) MIR-selected TDEs. These spectra reveal compact, high-ionization emission lines associated with SMBH accretion and strong silicate emission features indicative of optically thin dust emission. Coupled with weak emission from PAHs, the high-IP lines resemble MIR spectra of AGN, rather than star-forming galaxies that show strong PAH features and weak high-IP lines. The similarity to AGN MIR spectra is not particularly surprising, given that both TDEs and AGN are powered by SMBH accretion; the dust around these two systems is heated with similar ionizing radiation that destroys most PAHs and excites atomic gas. However, most AGN show relatively weak silicate features compared to the strong silicate emission features presented in this work. We suspect that this is a direct result of the different dust distributions around active and dormant SMBHs. The canonical clumpy, optically thick torus in AGN is thought to require radiation pressure support from a long-lived accretion flow (Pier & Krolik, 1992; Thompson et al., 2005; Krolik, 2007) that does not exist in TDEs. Thus, together with the accretion-driven emission lines, this strengthens the argument that these sources are indeed TDEs.
@@
Motivated by the strong silicate emission features, we have modeled the JWST MIRI spectra with a simple, optically thin dust model that accounts for the time-dependent spectral evolution. This model can accurately reproduce the strong silicate features and the long wavelength continuum emission. However, the short wavelength emission is dominated by dust near the SMBH heated by the late-time emission from the TDE. Since the initial flare is quite luminous, most of the small grains and silicates at such small radii will have been destroyed. Figure 7 shows that dust emission from larger ($a=0.5$ µm) graphite grains heated by a late-time plateau of $L=2\times 10^{42}$ erg s-1 can reproduce the near-IR excess at $\lambda\lesssim 8$ µm. Interestingly, near-IR excesses are also seen in AGN and have been similarly explained with hot ($T\gtrsim 1200$ K), pure graphite dust that can exist closer to the SMBH than silicate dust (e.g. Mor et al., 2009; Mor & Netzer, 2012). As TDEs may emit a sizable fraction of their total energy at late times in a viscously spreading accretion disk (Mummery & Balbus, 2020; Mummery, 2021; Mockler & Ramirez-Ruiz, 2021; Mummery & van Velzen, 2024), accurate estimation of the late-time plateaus and by association the close dust to the SMBH is crucial to our understanding of TDE energetics and solving the TDE missing energy problem (e.g. Lu & Kumar, 2018). We stress, however, that a more exhaustive search of this complex parameter space is necessary to assess the nature and strength of the late-time plateau in these IR-selected TDEs.
[…]
@@
Given the complexity associated with astrophysical dust in extreme environments, our modeling has made a number of key assumptions and approximations. One important assumption that we have made is that the dust is optically thin to its own emission (i.e. $\tau_{9.7\,\mu\mathrm{m}}\lesssim 1$). This assumption allows us to neglect scattering in our calculations, as a proper treatment of scattering is difficult to do in a non-spherical emitting surface given geometric effects. Our order of magnitude estimates suggest that optically thin dust is relevant for the timescales (i.e. distances) we are probing with JWST MIRI, but there may be dense, optically thick dust closer to the SMBH. Such dust would impact our calculations by attenuating the dominant UV emission and leading to the more dust being heated by the reprocessed radiation from the inner dust, rather than the intrinsic flare. Future efforts with Monte Carlo radiation transport simulations are necessary to assess the impact of scattering and heating from reprocessed emission on with the necessary time-dependent effects that exist in TDEs. Such simulations are now starting to be viable computationally, but to-date have only been done with gray opacities to estimate the light curve evolution (e.g. Tuna et al., 2025).
These JWST spectra have proven invaluable for assessing the composition and optical depth of the dust in the nuclei of TDE hosts. However, most studies of IR-bright TDEs rely solely on the two-band photometry from NEOWISE, and thus, it is prudent to compare the JWST spectra to the inferred MIR emission from NEOWISE photometry. Fitting the NEOWISE data with either a single temperate blackbody or modified blackbody systematically underpredicts the long wavelength emission in the JWST band. This difference is likely as a result of the multi-temperature nature of the circumnuclear environment; the NEOWISE data is predominantly probing the hotter late-time plateau at short wavelengths, while the JWST spectra are more sensitive to the further, cooler dust. This necessitates further exploration of the IR spectral evolution to facilitate accurate estimate of TDE energetics, which we plan to explore in a future publication by incorporating the NEOWISE evolution into this optically thin dust model.
@@
This modeling makes two key predictions that are relevant for next-generation IR observatories. At very late times ($\gtrsim 100$ years), the illumination of distant, cool dust will lead to far-IR dust echoes that may be detectable with next-generation far-IR observatories like the PRobe far-Infrared Mission for Astrophysics (PRIMA; Moullet et al., 2023). At very early times ($\lesssim 0.5$ years), the spectrum should be dominated by pure graphite dust, as all of the silicates near the SMBH will be destroyed. JWST observations in the first 6 months after a TDE are necessary to test this hypothesis and inform our modeling of the dust close to the SMBH. Although MIR-selected TDE candidate samples are contaminated by SNe at early times, recent work has shown that only a small fraction ($\sim 0.1\%$) of the dominant contaminants, Type Ia SNe, are MIR-bright (Mo et al., 2025), thereby making these early-time searches more tractable. Additionally, in the coming years, numerous new observatories will revolutionize time-domain IR studies with greatly improved cadence, including the Nancy Grace Roman Space Telescope (Spergel et al., 2015), the Wide-field INfrared Transient ExploreR (WINTER; Lourie et al., 2020), the PRime-focus Infrared Microlensing Experiment (PRIME [^4] http://www-ir.ess.sci.osaka-u.ac.jp/prime/index.html), and Cryoscope (Kasliwal et al., 2025) in the near-IR and NEOSurveyor (Mainzer et al., 2023) in the mid-IR. Likewise, SPHEREx (Doré et al., 2014), which launched in March 2025, will produce all-sky near-to-mid-IR spectra every 6 months that can help identify IR-bright TDEs earlier in their evolution. Thus, we expect that catching TDE dust echoes in their early phases of evolution will be possible with the next generation missions. These observations will break degeneracies between dust composition and accretion luminosity, thereby allowing us to answer fundamental questions related to the overall impact of TDEs on black hole growth.
@@
We thank the anonymous referee for their helpful comments. We thank Hagai Netzer and Andy Mummery for insightful discussions. This work is based on observations made with the NASA/ESA/CSA JWST. The data were obtained from the Mikulski Archive for Space Telescopes at the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-03127 for JWST. These observations are associated with program #3696. Support for program #3696 was provided by NASA through a grant from the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-03127. We acknowledge the support of the National Aeronautics and Space Administration through ADAP grant number 80NSSC24K0663. MG is supported in part by NASA XMM-Newton grant 80NSSC24K1885. The data presented in this paper were obtained from the Mikulski Archive for Space Telescopes (MAST) at the Space Telescope Science Institute. The specific observations analyzed can be accessed via https://doi.org/10.17909/efbj-mf87 (catalog https://doi.org/10.17909/efbj-mf87). STScI is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS5–26555. Support to MAST for these data is provided by the NASA Office of Space Science via grant NAG5–7584 and by other grants and contracts.